Blue Corporation began operations on December 1, 2019. The only inventory transaction in 2019 was the purchase of inventory on December 10, 2019, at a cost of $25 per unit. None of this inventory was sold in 2019. Relevant information is as follows.

Ending inventory units
   December 31, 2019 122
   December 31, 2020, by purchase date
      December 2, 2020 122
      July 20, 2020 50 172


During the year 2020, the following purchases and sales were made.

Purchases

Sales

March 15 322 units at $30 April 10 222
July 20 322 units at 32 August 20 322
September 4 222 units at 35 November 18 172
December 2 122 units at 38 December 12 222


The company uses the periodic inventory method.

Incorrect answer iconYour answer is incorrect.

Calculate average-cost per unit. (Round answer to 2 decimal places, e.e. 2.76.)

Average-cost

$

eTextbook and Media

Incorrect answer iconYour answer is incorrect.

Determine ending inventory under (1) specific identification, (2) FIFO, (3) LIFO, and (4) average-cost. (Round answers to 0 decimal places, e.g. 2,760.)

Specific Identification

FIFO

LIFO

Average-Cost

Ending Inventory

$

$

$

$

eTextbook and Media

Incorrect answer iconYour answer is incorrect.

Calculate price index. (Round answer to 4 decimal places, e.g. 2.7653.)

Price Index

eTextbook and Media

Incorrect answer iconYour answer is incorrect.

Determine ending inventory using dollar-value LIFO. Assume that the December 2, 2017, purchase cost is the current cost of inventory. (Hint: The beginning inventory is the base layer priced at $25 per unit.) (Round answer to 0 decimal places, e.g. 2,760.)

Ending inventory at dollar-value LIFO

$
Answer #1

Average Cost Per Unit

Weighted Average Unit Cost = Total Cost of Inventory
Total Units in Inventory
Units Unit Cost ($ per unit) Total Cost
January 1, 2020 Beginning Inventory 122 25 3050
Purchases in 2020:
Mar-15 322 30 9660
Jul-20 322 32 10304
Sep-04 222 35 7770
Dec-02 122 38 4636
1110 31.91 35420
(*) 35420/1110

Ending Inventory:

1. Specific Identification:

As per Specific Identification method, Ending Inventory is made up of 122 units from Dec 2, 2020 purchases and 50 units from July 20, 2020 purchases
Ending Inventory ($) = (122 * 38) + (50 * 32)

= 6236

2. FIFO

As per FIFO, units purchased first are sold first. Hence Ending inventory which is 172 units are left from 122 units @ $38/ unit purchased on Dec 02, 2020 and 50 units @ $35/ unit purchased on Sep 04, 2020
Ending Inventory ($) = (122 * 38) + (50 * 35)
= 6386

3. LIFO

As per LIFO, units purchased Last are sold first. Hence Ending inventory which is 172 units are left from 122 units @ $25/ unit purchased on Dec 10, 2019 and 50 units @ $30/ unit purchased on Mar 15, 2020
Ending Inventory ($) = (122 * 25) + (50 * 30)
= 4550

4. Average Cost

Units Available for sale (A) = 122 + 322 + 322 + 222 + 122 1110
Units Sold (B) = 222 + 322 + 172 + 222 938
Units in Ending Inventory (A- B) 172
Weighted Average Unit Cost ($) = 35420
1110
= 32
Ending Inventory ($) = 32 * 172
= 5504